An individual plots her utility from a consumption bundle based on the three utility structures as follows Linear: U(x y) 5x+3y Cobb-Douglas: U(x,y)-хгуч CES: (3x2 + (-2)y2)i2 The unit price of x and y are 5 and 2 respectively. 212-1 a. Compute the utility for different structures for the following bundles - (5,5), (1,8) and (25,5) b. Compute the marginal rate of substitution for the three utility structures c. Given a budget of 100 units, determine the optimal allocation for each of the three utility structures
Answer #1

An individual plot her utility from a consumption bundle, he has three utility structures a) Utility for different structuresU(5,5)=(3x52-2x52) 2, U(18) (3x12 -2x2 -(-125F(maginery) U (x,y)- (3x2 -2y) at (25,5) U(25,5) (3x252 -2x5*yu2y = 42.72 b) CoMRS MU Br6x c) The price of x and y are 5 and 2 respectively Given a budget of 100 units The budget line is 5x+2y-100 Optimal-2y 5 5x +2y= 100 5x1 그1+2),-100 -4y+2y = 100 41 y =-50 -59 4x (-50) For CESU(x,y)-(3r-2y2) 2) MRS = 2), 2 -6x = 103. 51+21,